OpenStudy (anonymous):

Write out the formulas and show your work. You're given a sample with n=8 measurements: 3, 1, 5, 6, 4, 4, 3,5 A. Calculate the range (maximum value ­ minimum value). B. Calculate the sample mean. C. Calculate the sample variance and sample standard deviation. D. The range is approximately how many sample standard deviations?

OpenStudy (unklerhaukus):

so what are the max and min values in that set?

OpenStudy (anonymous):

is it 1 and 5?

OpenStudy (anonymous):

i mean 6

OpenStudy (unklerhaukus):

1 is min and 6 is the max , right

OpenStudy (unklerhaukus):

so the range is ...

OpenStudy (anonymous):

It seems your home work

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

@Rohangrr

OpenStudy (anonymous):

Oh, I'm sorry, I get the first part and second part, just not the standard deviation and sample variance part and the question under it

OpenStudy (unklerhaukus):

yeah for A the range is is 5 what did you get for B

OpenStudy (anonymous):

3.875?

OpenStudy (anonymous):

is the mean

OpenStudy (anonymous):

@rohangrr

OpenStudy (unklerhaukus):

good work

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

hello? still with me?

OpenStudy (unklerhaukus):

\[s^2=\frac{1}{n-1}\sum(X-\overline X)^2\]

OpenStudy (anonymous):

Is the formula for standard deviation?

OpenStudy (unklerhaukus):

s is standard deviation (for a sample)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

Can you please guide me through how to complete it and what the symbols mean?

OpenStudy (unklerhaukus):

i can do standard deviation by i get too confused when it is for a sample only i used \(\overline X\) for average

OpenStudy (anonymous):

okay and average would be the mean?

OpenStudy (anonymous):

I'm sorry I really don't understand this still.. haha

OpenStudy (anonymous):

I think this is the equation, but I don't get it....

OpenStudy (unklerhaukus):

yeah average = mean